# Taming VR Teleoperation and Learning from Demonstration for Multi-Task Bimanual Table Service Manipulation

### v1

@@ Abstract
This technical report presents the champion solution of the Table Service Track in the ICRA 2025 What Bimanuals Can Do (WBCD) competition. We tackled a series of demanding tasks under strict requirements for speed, precision, and reliability: unfolding a tablecloth (deformable-object manipulation), placing a pizza onto the table (pick-and-place), and opening and closing a food container with the lid. Our solution combines VR-based teleoperation and Learning from Demonstrations (LfD) to balance robustness and autonomy. Most subtasks were executed through high-fidelity remote teleoperation, while the pizza placement was handled by an ACT-based policy trained from 100 in-person teleoperated demonstrations with randomized initial configurations. By carefully integrating scoring rules, task characteristics, and current technical capabilities, our approach achieved both high efficiency and reliability, ultimately securing the first place in the competition.
@@ Acknowledgements
We would like to thank the organizers of the ICRA 2025 WBCD Challenge: Xu Zhuo, Dennis Gu and Peter Yu, for their on-site support in coordinating hardware materials and clarifying competition details for our team. We also thank the final-round judges, Di Huang and Yuanzhe Dong, for their efforts. Our appreciation extends to Josh Zhang, Nina Ma and Yongzhen Zou from ARX Robotics for their hardware support during the competition and for sponsoring the championship prize.

### v2

@@ Abstract
This technical report presents the champion solution of the Table Service Track in the ICRA 2025 What Bimanuals Can Do (WBCD) competition. We tackled a series of demanding tasks under strict requirements for speed, precision, and reliability: unfolding a tablecloth (deformable-object manipulation), placing a pizza into the container (pick-and-place), and opening and closing a food container with the lid. Our solution combines VR-based teleoperation and Learning from Demonstrations (LfD) to balance robustness and autonomy. Most subtasks were executed through high-fidelity remote teleoperation, while the pizza placement was handled by an ACT-based policy trained from 100 in-person teleoperated demonstrations with randomized initial configurations. By carefully integrating scoring rules, task characteristics, and current technical capabilities, our approach achieved both high efficiency and reliability, ultimately securing the first place in the competition.
@@ Acknowledgements
We would like to thank the organizers of the ICRA 2025 WBCD Challenge: Zhuo Xu, Dennis Gu and Peter Yu, for their on-site support in coordinating hardware materials and clarifying competition details for our team. We also thank the final-round judges, Di Huang and Yuanzhe Dong, for their efforts. Our appreciation extends to Josh Zhang, Nina Ma and Yongzhen Zou from ARX Robotics for their hardware support during the competition and for sponsoring the championship prize.